# A Scalable MCMC Algorithm for Bayesian Inference on Binary Model Spaces

### 27 Mar 2026 v2

@@ 3.3 Additional discussion and results
At every iteration $s$, our DTMC can jump to any state $m^{\prime}$ with positive posterior probability, since $P_{\varepsilon_{s}}(m,m^{\prime})>0$ whenever $\pi(m^{\prime})>0$. This comes with the evident advantage of excellent mixing. However, it also yields a noteworthy observation: our DTMC can visit states $m^{\prime}$ with very low posterior probability, particularly for larger $\varepsilon_{s}$. Imagine for example the simple two-dimensional model space $\mathcal{M}=\{0,1\}^{2}$ and suppose that the posterior probabilities are $\pi(1,1)=0.01$, and $\pi(0,0)=\pi(0,1)=\pi(1,0)=0.33$. An MCMC algorithm targeting $\pi$ should therefore rarely visit the state $(1,1)$. For example, the transition rates to $(1,1)$ of the BD process in Lemma 2.1 are $Q((0,0),(1,1))=0$ and $Q((0,1),(1,1))=Q((1,0),(1,1))=0.03$. Our DTMC, however, is more likely to visit such a state, especially for higher $\varepsilon_{s}$. For example, when in state $(0,0)$ and when $\varepsilon_{s}=0.9$, our DTMC has a chance of $P_{\varepsilon_{s}}((0,0),(1,1))=\varepsilon_{s}^{2}=0.81$ of moving from $(0,0)$ to $(1,1)$. At first glance, this might seem problematic. However, note that Theorem 3.1 suggests a small value for the constant $\varepsilon$ and Theorem 3.2 requires $\lim_{s\ \to\infty}\varepsilon_{s}=0$. These conditions ensure that, in the long run, jumps to states with low posterior probabilities become infrequent, and that the DTMC’s limiting distribution will get arbitrarily close to that of the BD process.
@@
or rejected otherwise. While the accept/reject step renders the algorithm exact, it slows it down, as a drawback.
@@
Lemma $3.$
Consider a binary model space $\{0,1\}^{k}$ and a factorizable posterior $p(m\mid y)$, as in (10). Consider a BD process as in Lemma 2.1 with rates $q_{i}(m)$. Moreover, consider a homogeneous, discrete-time MCMC with transition probabilities (8), with $\varepsilon_{s}=\varepsilon\in(0,1)$ for all $s$, and where $q_{i}(m)$ are the rates of the BD process. Then, for any $\varepsilon\in(0,1)$ the DTMC has the same invariant distribution as the BD process.
Lemma 3.3 provides intuition to the workings of Algorithm 1. The factorizable posterior ensures that the elements $m_{1},\cdots,m_{k}$ are independent. This element-wise posterior independence is exactly why the homogeneous variant of Algorithm 1 works for all $\varepsilon\in(0,1)$, since Algorithm 1 updates the elements $m_{i}$ independent of each other. In most cases, however, posteriors are not factorizable. That is why, in the general case, we need either the condition $\varepsilon\to 0$ (Theorem 3.1) or both decay and divergence conditions (Theorem 3.2).
Extension to joint posterior: So far, we have presented an extension to the BD process (Lemma 2.1) that creates a superior DTMC (Algorithm 1). Both the BD process and our DTMC are suitable for inference on the model posterior (2), but not to learn the parameters $\theta_{m}$ associated with each model. That is, the algorithms are not immediately applicable for inference on the joint posterior (1). The step to joint inference, however, is not large. In fact, the next lemma presents a straightforward strategy to use the samples $(m^{(1)},\ldots,m^{(S)})$ from Algorithm 1 to obtain samples $((m^{(1)},\theta^{(1)}_{m}),\ldots,(m^{(S)},\theta^{(S)}_{m})$ from the joint posterior (1).
@@
Let $m^{(1)},m^{(2)},\ldots$ be the states of a discrete-time Markov chain with limiting distribution $\pi(m)$. If we now sample the parameter vectors $\theta^{(1)},\theta^{(2)},\ldots$ from a distribution $\pi(\theta\mid m^{(s)})$ for all $s=1,2,..$. Then, the resulting samples $(m^{(1)},\theta^{(1)}),(m^{(2)},\theta^{(2)}),\ldots$ are samples from a Markov chain with limiting distribution $\pi(m)\pi(\theta\mid m)$
@@ 4 Applications
Algorithm 1 can be used for Bayesian model inference on problems with a binary model space. This section presents a well-known example: undirected Gaussian graphical models. In Supplementary Material B, we also apply our approach to Bayesian variable selection and Ising models. For each application, the path to construct our Multiple Jump algorithm is the same: we formulate the model posterior (2) and specify the rates $q_{i}(m):=Q(m,m^{i})$ for the BD process in Lemma 2.1. Using these rates in Algorithm 1 will give the desired algorithm.
4.1 Undirected Gaussian graphical models
This subsection treats undirected Gaussian graphical models (GGMs). In fact, the BD process in Lemma 2.1 is already proposed by Mohammadi et al. (2024). This subsection reviews those existing results and presents our algorithm as an extension of this BD process.
@@ 4.1 Undirected Gaussian graphical models
GGMs assume that the data $y$ come from a multivariate normal distribution. More precisely, $y$ is an $n\times p$ matrix containing $n$ observations of the variables $Y_{1},\ldots,Y_{p}$. These variables follow a multivariate normal distribution with unknown $p\times p$ covariance matrix $\Sigma$ and unknown precision matrix $K:=\Sigma^{-1}$. The matrix $K$, with elements $k_{ij}$, encodes the conditional dependency between any two variables $Y_{i}$ and $Y_{j}$, because
@@
If we denote the presence of an edge with a 11 and the absence of an edge with a 0, we can encode every graph as binary vector of length k=p(p1)/2k=p(p-1)/2. The graphs $G$ are therefore binary “models”, and the space of all $p\times p$ graphs is a binary model space. The precision matrix $K$ is the model-associated parameter. Given a graph $G$, the prior on the precision matrix is denoted by $p(K\mid G)$. Popular choices include the G-wishart prior (Roverato, 2002) or the spike-and-slab prior (Wang, 2015). See Vogels et al. (2024) for a comprehensive review on Bayesian structure learning in GGMs.
@@
where $y_{h}$ is a vector containing the $n$ observations of variable $h$, $nb(h,G)$ refers to the set of neighbors of node $h$ with respect to $G$ and $y_{A}$ is the sub-matrix of $y$ corresponding to the nodes that are in a set AA. This approximation is remarkably fast, because it only requires the evaluation of the four terms $p(y_{h}\mid y_{nb(h,G)})$, each of which can be obtained quickly using another approximation (Leppä-aho et al., 2017)[Equation 9]. Mohammadi et al. (2024) use equation (11) to design a BD process as described in Lemma 2.1 with rates
@@
$q_{e}(G):=\min\left(1,\frac{\hat{p}(y\mid G^{e})}{\hat{p}(y\mid G)}\frac{p(G^{e})}{p(G)}\right)$ (12)
The algorithm of Mohammadi et al. (2024) outperforms other state-of-the art algorithms, but still takes more than an hour to solve $1000$-variable problems (see Section 5.1), mostly because the algorithm – being a BD process – updates only one edge per iteration.
@@
Our work allows us to extend the BD process of Mohammadi et al. (2024) to a DTMC that does allow multiple edge updates per iteration. Our DTMC will have transition probabilities (8) with $q_{e}(G)$ as in (12). By Theorem 3.2, the DTMC has the same limiting distribution as the BD process, provided that the sequence $\{\varepsilon_{s}\}_{s\geq 1}$ satisfies (9). Moreover, due to Lemma 3.4, if we sample precision matrices $K^{(s)}$ from $p(K\mid G^{(s)},y)$ for $s=1,\ldots,S$, the resulting samples $(G^{(1)},K^{(1)}),\ldots,(G^{(S)},K^{(S)})$ are samples from the joint posterior $p(K,G\mid y)$. We tested our approach in a simulation study (Section 5.1) and on real-life data (Section 5.2).
@@ 5.1 Simulation study
This section highlights the scalability of our approach in a simulation setting. We compare our approach (Algorithm 1) to the state-of-the-art algorithm in Bayesian structure learning in Gaussian graphical models: the BD-MPL algorithm (Mohammadi et al., 2024), which is up to $10$ times faster than the competing algorithms on large-scale instances ($p=1000$). We test our algorithm with homogeneous transition probabilities, using $\varepsilon\in\{0.001,0.01,0.3,0.7,0.95\}$, and inhomogeneous transition probabilities (8) where we use the slowly and quickly decaying sequences
$\varepsilon_{s}=0.3\frac{1}{\log_{10}(s+9)}\text{ for }s=1,\ldots,500\quad\text{ and }\quad\varepsilon_{s}=0.3\left(\frac{1}{s\log_{2}(s+1)}\right)^{0.4}\text{ for }s=1,\ldots,40000,$ (13)
respectively. Figure 1 shows both sequences. We report our results on an instance with $p=1000$ variables, $n=400$ observations, and edge density of $0.2\%$. This edge density might seem low, but note that under this density nodes are connected to two other nodes on average. The experiments were repeated eight times. Other simulation settings are described in Supplementary Material C.
@@
Figure 2 contains the results. It leads to three main conclusions: First, our approach is remarkably fast. In the homogeneous case for $\varepsilon_{s}\in\{0.3,0.7,0.95\}$, our approach solves $1000$-variable problems in under $30$ seconds. That is around $100$ to $200$ times faster than the state-of-the-art BD-MPL algorithm. In the inhomogeneous case, the quickly and the slowly decaying sequences are even faster. Second, for small values of $\varepsilon$, i.e. $\varepsilon\in\{0.001,0.01\}$ our approach slows down and starts behaving like the BD-MPL algorithm. This is in line with Theorem 3.1. Third, higher values of $\varepsilon$ do not lead to worse accuracy. On the contrary, all values of $\varepsilon$ as well as the quickly and the slowly decaying sequences lead to similar AUC-PR values. Lemma 3.3 might provide a reason for this surprising result. Using the G-wishart prior, the resulting posterior is not factorizable as in (10). However, for large values of $p$ and for sparse instances, we hypothesize that it does behave approximately factorizable. In line with Lemma 3.3, this ensures that the limiting distribution of Algorithm 1 is similar to that of the BD-MPL approach for all values of $\varepsilon$. This hypothesis is supported by Figures S9-S16 in Supplemenary Material C.3 and C.4. These figures show that, in denser cases, high values of $\varepsilon$ (e.g. $\varepsilon=0.75$) do in fact lead to worse accuracy.
@@ 5.2 Real-life data
The simulation study shows that our approach offers a significant reduction in computation time compared to the state-of-the-art. This section will show that this also holds for real-life data. Specifically, we will test Algorithm 1 on a dataset with $p=623$ variables and $n=653$ observations, and show that we obtain the same quality of inference as the state-of-the-art MPL-BD algorithm (Mohammadi et al., 2024), but up to $10$ times as fast.
The data set in question is the GSE15907 microarray dataset from Painter et al. (2011) and Desch et al. (2011), comprising gene expression data from 24,922 genes in 653 mouse immune cells, obtained from the Immunological Genome Project (Heng et al., 2008). Studying the conditional dependency structure among genes is important as it hints at how genes in immune cells work together to produce proteins impacting the body’s response to diseases. Understanding these mechanisms is essential to improve disease treatment strategies. The dataset was used in the context of Bayesian inference in GGMs before by both Chandra et al. (2024) and Mohammadi et al. (2024).
@@
We use the Bernoulli prior for the graph with a prior edge inclusion probability of $0.01$, based on the results of Chandra et al. (2024). We will run Algorithm 1 with homogeneous transition probabilities $\varepsilon\in\{0.001,0.01,0.3,0.6,0.9\}$ and inhomogeneous transition probabilities, where we use the slowly and quickly decaying sequences (13). We compare the performance of our algorithm with the MPL-BD algorithm by Mohammadi et al. (2024).
@@ 6 Conclusion
This work introduces an algorithm for Bayesian model inference on binary model spaces. We take a BD process as a starting point and show that a simple extension creates a new DTMC. Under conditions this DTMC has the same limiting distribution as the BD process. However, it can cross the entire model space in a single iteration, leading to remarkable computational efficiency. When applied to Gaussian graphical models, our algorithm is up to $200$ times as fast as the state-of-the art. We show that our framework can be applied to Ising models and variable selection too. There are several ways in which future work can build on this article.
First, more simulations are needed to highlight the strengths and weaknesses of our approach. We showed how our algorithm performs in a simulation study and on real-life data. However, it remains unclear how our algorithm performs on instances where $n\gg p$, instances with more than $1,000$ variables or very dense instances. Moreover, Supplementary Material B.1 and B.2 promise a reduction in computation time for variable selection and Ising models, respectively. Simulation studies are needed to check whether this promise can be fulfilled.
Secondly, our framework can be applied beyond the examples given in this paper. Starting with other binary models, such as mixed graphical models (Florez et al., 2025), decision trees (Mohammadi et al., 2020), or directed graphical models. Moreover, every discrete model space can be written as a binary model space with a constraint. This opens up the possibility to apply our work to Bayesian inference on discrete model spaces. Lastly, Chen (2023)[Section 4.4] show that an algorithm designed for a discrete state space can be extended to work on a continuous state space. A similar extension could apply our framework to continuous model spaces.
Thirdly, our framework can be extended for inference on the joint posterior. Algorithm 1 focuses on inference on the model posterior (2). There are two ways to realize this extension. First, we can follow Lemma 3.4 and sample the model-associated parameters after running Algorithm 1. Simulations are needed to show whether this works in practice. Alternatively, we can take a BD process on the joint space as a starting point (e.g. Mohammadi and Wit (2015)), and tweak Algorithm 1 to extend this to a superior DTMC on the joint space. A theoretical analysis and simulation study is needed to assess this strategy.
Fourthly, further reductions in computational time are possible. One such strategy could be to update only a subset of the rates at each iteration. Chen (2023)[Section 4.2] proposes this strategy in a similar set-up, and shows its potential to reduce computational costs. Alternatively, we can achieve higher efficiency by calculating the rates using only a subset of the observations, as in Welling and Teh (2011).

### 02 Apr 2026 v3

@@ 3.3 Additional discussion and results
At every iteration $s$, our MJ-MCMC algorithm can jump to any state $m^{\prime}$ with positive posterior probability, since $P_{\varepsilon_{s}}(m,m^{\prime})>0$ whenever $\pi(m^{\prime})>0$. This comes with the evident advantage of excellent mixing. However, it also yields a noteworthy observation: our DTMC can visit states $m^{\prime}$ with very low posterior probability, particularly for larger $\varepsilon_{s}$. Imagine for example the simple two-dimensional model space $\mathcal{M}=\{0,1\}^{2}$ and suppose that the posterior probabilities are $\pi(1,1)=0.01$, and $\pi(0,0)=\pi(0,1)=\pi(1,0)=0.33$. An MCMC algorithm targeting $\pi$ should therefore rarely visit the state $(1,1)$. For example, the transition rates to $(1,1)$ of the BD process in Lemma 2.1 are $Q((0,0),(1,1))=0$ and $Q((0,1),(1,1))=Q((1,0),(1,1))=0.03$. Our DTMC, however, is more likely to visit such a state, especially for higher $\varepsilon_{s}$. For example, when in state $(0,0)$ and when $\varepsilon_{s}=0.9$, our DTMC has a chance of $P_{\varepsilon_{s}}((0,0),(1,1))=\varepsilon_{s}^{2}=0.81$ of moving from $(0,0)$ to $(1,1)$. At first glance, this might seem problematic. However, note that Theorem 3.1 suggests a small value for the constant $\varepsilon$ and Theorem 3.2 requires $\lim_{s\ \to\infty}\varepsilon_{s}=0$. These conditions ensure that, in the long run, jumps to states with low posterior probabilities become infrequent, and that the DTMC’s limiting distribution will get arbitrarily close to that of the BD process.
@@
or rejected otherwise. The accept/reject step renders the algorithm exact, but requires extra computational cost.
@@
Consider a binary model space $\{0,1\}^{k}$ and a factorizable posterior $p(m\mid y)$, as in (10). Consider a BD process as in Lemma 2.1 with rates $q_{i}(m)$. Moreover, consider a homogeneous, discrete-time MCMC with transition probabilities (8), with $\varepsilon_{s}=\varepsilon\in(0,1)$ for all $s$, and where $q_{i}(m)$ are the rates of the BD process. Then, for any $\varepsilon\in(0,1)$, the DTMC has the same invariant distribution as the BD process.
Using a BD process or Algorithm 1 to recover a factorizable posterior (10) is a form of over-engineering, since one can instead decompose the posterior into $k$ element-wise posteriors $p(m_{i}\mid y)$, each of which can be obtained easily using a two-state ($m_{i}=0$ or $m_{i}=1$) Markov chain. Lemma 3.3 is therefore not meant as a proposal for a fast algorithm, but instead to provide intuition to the workings of Algorithm 1. The factorizable posterior ensures that the elements $m_{1},\cdots,m_{k}$ are independent. This element-wise posterior independence is exactly why the homogeneous variant of Algorithm 1 works for all $\varepsilon\in(0,1)$, since Algorithm 1 updates the elements $m_{i}$ independent of each other. In most cases, however, posteriors are not factorizable. That is why, in the general case, we need either the condition $\varepsilon\to 0$ (Theorem 3.1) or both decay and divergence conditions (Theorem 3.2).
Extension to joint posterior: So far, we have presented an extension to the BD process (Lemma 2.1) that creates a superior DTMC (Algorithm 1). Both the BD process and our DTMC are suitable for inference on the model posterior (2), but not to learn the parameters $\theta_{m}$ associated with each model. That is, the algorithms are not immediately applicable for inference on the joint posterior (1). The step to joint inference, however, is straightforward. In fact, the next lemma presents a simple strategy to use the samples $(m^{(1)},\ldots,m^{(S)})$ from Algorithm 1 to obtain samples $((m^{(1)},\theta^{(1)}_{m}),\ldots,(m^{(S)},\theta^{(S)}_{m})$ from the joint posterior (1).
Lemma 3.4.
Let $m^{(1)},m^{(2)},\ldots$ be the states of a discrete-time Markov chain with limiting distribution $\pi(m)$. If we now sample the parameter vectors $\theta^{(1)},\theta^{(2)},\ldots$ from a distribution $\pi(\theta\mid m^{(s)})$ for all $s=1,2,..$. Then, the resulting samples $(m^{(1)},\theta^{(1)}),(m^{(2)},\theta^{(2)}),\ldots$ are samples from a Markov chain with limiting distribution $\pi(m)\pi(\theta\mid m)$.
@@ 4 Applications
The MJ-MCMC algorithm (Algorithm 1) can be used for Bayesian model inference on problems with a binary model space. In Supplementary Material B, we apply our approach to Bayesian variable selection and Ising models. In this section, we present a different example: undirected Gaussian graphical models (GGMs). GGMs are a good starting point, because it is a relatively well-known field. In fact, the BD process in Lemma 2.1 is already proposed by Mohammadi et al. (2024). This subsection reviews those existing results and presents our algorithm as an extension of this BD process.
@@
GGMs assume that the data $y$ come from a multivariate normal distribution (Lauritzen, 1996). More precisely, $y$ is an $n\times p$ matrix containing $n$ observations of the variables $Y_{1},\ldots,Y_{p}$. These variables follow a multivariate normal distribution with unknown $p\times p$ covariance matrix $\Sigma$ and unknown precision matrix $K:=\Sigma^{-1}$. The matrix $K$, with elements $k_{ij}$, encodes the conditional dependency between any two variables $Y_{i}$ and $Y_{j}$, because
@@
If we denote the presence of an edge with a $1$ and the absence of an edge with a 0, we can encode every graph as a binary vector of length k=p(p1)/2k=p(p-1)/2. The graphs $G$ are therefore binary “models”, and the space of all $p\times p$ graphs is a binary model space. The precision matrix $K$ is the model-associated parameter. Given a graph $G$, the prior on the precision matrix is denoted by $p(K\mid G)$. Popular choices include the G-wishart prior (Roverato, 2002) or the spike-and-slab prior (Wang, 2015). See Vogels et al. (2024) for a comprehensive review on Bayesian structure learning in GGMs.
@@
where $y_{h}$ is a vector containing the $n$ observations of variable $h$, $nb(h,G)$ refers to the set of neighbors of node $h$ with respect to $G$ and $y_{A}$ is the sub-matrix of $y$ corresponding to the nodes that are in a set $A$. This approximation is remarkably cheap to evaluate, because it only requires the evaluation of the four terms $p(y_{h}\mid y_{nb(h,G)})$, each of which can be obtained quickly using another approximation (Leppä-aho et al., 2017)[Equation 9]. Mohammadi et al. (2024) use equation (11) to design a BD process as described in Lemma 2.1 with rates
@@
$q_{e}(G):=\min\left\{1,\frac{\hat{p}(y\mid G^{e})}{\hat{p}(y\mid G)}\frac{p(G^{e})}{p(G)}\right\}.$ (12)
The algorithm of Mohammadi et al. (2024) outperforms other state-of-the-art algorithms, but still takes more than an hour to solve $1000$-variable problems (see Section 5.1), mostly because the algorithm – being a BD process – updates only one edge per iteration.
@@
Our work allows us to extend the BD process of Mohammadi et al. (2024) to a DTMC that does allow multiple edge updates per iteration. Our DTMC will have transition probabilities (8) with $q_{i}(m):=q_{e}(G)$ as in (12). By Theorem 3.2, the DTMC has the same limiting distribution as the BD process, provided that the sequence $\{\varepsilon_{s}\}_{s\geq 1}$ satisfies (9). Moreover, similar to (Mohammadi et al., 2024)[Sectiopn 4.3], we can sample precision matrices $K^{(s)}$ from $p(K\mid G^{(s)},y)$ for $s=1,\ldots,S$. Due to Lemma 3.4, the resulting samples $(G^{(1)},K^{(1)}),\ldots,(G^{(S)},K^{(S)})$ are then samples from the joint posterior $p(K,G\mid y)$. We tested our approach in a simulation study (Section 5.1) and on real-life data (Section 5.2).
@@ 5.1 Simulation study
This section highlights the scalability of our approach in a simulation setting. We compare our approach (Algorithm 1) to the state-of-the-art algorithm in Bayesian structure learning in Gaussian graphical models: the BD-MPL algorithm (Mohammadi et al., 2024), which is up to $10$ times faster than the competing algorithms on large-scale instances ($p=1000$) (Vogels et al., 2024). We test our algorithm with homogeneous transition probabilities, using $\varepsilon\in\{0.001,0.01,0.3,0.7,0.95\}$, and inhomogeneous transition probabilities (8) where we use the slowly and quickly decaying sequences
$\varepsilon_{s}=0.3\frac{1}{\log_{10}(s+9)}\quad\text{ and }\quad\varepsilon_{s}=0.3\left(\frac{1}{s\log_{2}(s+1)}\right)^{0.4}\text{ for }s=1,2\ldots$ (13)
respectively. Figure 1 shows both sequences. We report our results on an instance with $p=1000$ variables, $n=400$ observations, and edge density of $0.2\%$. This edge density might seem low, but note that under this density, nodes are connected to two other nodes on average. The experiments were repeated eight times. Other simulation settings are described in Supplementary Material C.
@@
Figure 2 contains the results. It leads to three main conclusions: First, our approach is remarkably fast. In the homogeneous case for $\varepsilon_{s}\in\{0.3,0.7,0.95\}$, our approach solves $1000$-variable problems in under $30$ seconds. That is around $100$ to $200$ times faster than the state-of-the-art BD-MPL algorithm. In the inhomogeneous case, the quickly and the slowly decaying sequences are even faster. Second, for small values of $\varepsilon$, i.e., $\varepsilon\in\{0.001,0.01\}$, our approach slows down and starts behaving like the BD-MPL algorithm. This is in line with Theorem 3.1. Third, higher values of $\varepsilon$ do not lead to worse accuracy. On the contrary, all values of $\varepsilon$ as well as the quickly and the slowly decaying sequences lead to similar AUC-PR values. Lemma 3.3 might provide a reason for this surprising result. Using the G-wishart prior, the resulting posterior is not factorizable as in (10). However, for large values of $p$ and for sparse instances, we hypothesize that it does behave approximately factorizable. In line with Lemma 3.3, this ensures that the limiting distribution of Algorithm 1 is similar to that of the BD-MPL approach for all values of $\varepsilon$. This hypothesis is supported by Figures S9-S16 in Supplemenary Material C.3 and C.4. These figures show that, in denser cases, high values of $\varepsilon$ (e.g. $\varepsilon=0.75$) do in fact lead to worse accuracy.
@@ 5.2 Real-life data
The simulation study shows that our approach offers a significant reduction in computation time compared to the state-of-the-art. This section will show that this also holds for real-life data. Specifically, we will test Algorithm 1 on a dataset with $p=623$ variables and $n=653$ observations, and show that we obtain the same quality of inference as the state-of-the-art BD-MPL algorithm (Mohammadi et al., 2024), but up to $10$ times as fast.
The data set in question is the GSE15907 microarray dataset from Painter et al. (2011) and Desch et al. (2011), comprising gene expression data from 24,922 genes in 653 mouse immune cells, obtained from the Immunological Genome Project (Heng et al., 2008). Studying the conditional dependency structure among genes is important as it hints at how genes in immune cells work together to produce proteins impacting the body’s response to diseases. Understanding these mechanisms is essential to improving disease treatment strategies. The dataset was used in the context of Bayesian inference in GGMs before by both Chandra et al. (2024) and Mohammadi et al. (2024).
@@
We use the Bernoulli prior for the graph with a prior edge inclusion probability of $0.01$, based on the results of Chandra et al. (2024). We will run Algorithm 1 with homogeneous transition probabilities $\varepsilon\in\{0.001,0.01,0.3,0.6,0.9\}$ and inhomogeneous transition probabilities, where we use the slowly and quickly decaying sequences (13). We compare the performance of our algorithm with the BD-MPL algorithm by Mohammadi et al. (2024). We initialize both algorithms at the empty graph.
@@ 6 Conclusion
This work introduces an algorithm for Bayesian model inference on binary model spaces. We take a BD process as a starting point and show that a simple extension creates a new DTMC. Under mild conditions, this DTMC has the same limiting distribution as the BD process. However, it can cross the entire model space in a single iteration, leading to remarkable computational efficiency. When applied to Gaussian graphical models, our algorithm is up to $200$ times as fast as the state-of-the-art. We show that our framework can be applied to Ising models and variable selection, too. There are several ways in which future work can build on this article.
First, our framework can be applied beyond the examples given in this paper. Starting with other binary models, such as mixed graphical models (Florez et al., 2025), decision trees (Mohammadi et al., 2020), mixture distributions (Stephens, 2000) or directed graphical models (Kuipers et al., 2022). Moreover, every discrete model space can be written as a binary model space with a constraint. This opens up the possibility of applying our work to Bayesian inference on discrete model spaces. Lastly, Chen (2023)[Section 4.4] show that an algorithm designed for a discrete state space can be extended to work on a continuous state space. A similar extension could apply our framework to continuous model spaces.
Second, our framework can be extended for inference on the joint posterior. Algorithm 1 focuses on inference on the model posterior (2). There are two ways to realize this extension. First, we can follow Lemma 3.4 and sample the model-associated parameters after running Algorithm 1. Simulations are needed to show whether this works in practice. Alternatively, we can take a BD process on the joint space as a starting point (e.g. Mohammadi and Wit (2015)), and tweak Algorithm 1 to extend this to a superior DTMC on the joint space. A theoretical analysis and simulation study are needed to assess this strategy.
Third, further reductions in computational time are possible. One such strategy could be to update only a subset of the rates at each iteration. Chen (2023)[Section 4.2] proposes this strategy in a similar set-up, and shows its potential to reduce computational costs. Alternatively, we can achieve higher efficiency by calculating the rates using only a subset of the observations, as in Welling and Teh (2011).